Circular No. 7749

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2001 QW_{322}

J. J. Kavelaars, McMaster University; J.-M. Petit, Observatoire de Besançon; B. Gladman, Observatoire de la Côte d’Azur; and M. Holman, Harvard-Smithsonian Center for Astrophysics, report the discovery of another putative binary transneptunian system, with available astrometry, orbital elements, and an ephemeris given on MPEC 2001-V34. Images acquired on Aug. 24 by Kavelaars at the Canada-France-Hawaii Telescope and analyzed by Petit revealed the presence of two transneptunian objects with identical sky motions and nearly identical magnitudes (R = 24.0 ± 0.1). The system has an angular separation of 4′′, several times larger than that of the other known cubewano binary system (cf. IAUC 7610) and the two other known plutino binary systems (cf. IAUC 3241, 3509; IAUC 7733). A gradual separation of 0′′.2 ± 0′′.1 per month is detected at the limit of the astrometric precision. An assumed 4-percent albedo implies that both components have radii of ∼ 100 km; assuming a circular orbit and a density of 1 g/cm^{3} indicates an orbital period of ≈ 4 yr, consistent with the measured change in angular separation. The very small variation in position angle implies that the system is seen nearly edge-on.

U SCORPII B. E. Schaefer, University of Texas, Austin, writes: “An eruption of

the recurrent nova U Sco was discovered on 29 archival photographs at the Harvard College Observatory dating from mid-1945. The observed peak magnitude was at B = 9.57 on May 31, although the lack of plates over the previous 24 days allows for a somewhat brighter maximum. Representative B magnitudes: June 5, 11.56; 11, 13.74; July 2, 15.15; 7, 16.98; 16, 17.37. The 1945 event suggests that the average recurrence timescale is fairly uniform at around 11 years, with outbursts missed (perhaps due to conjunction with the sun) around 1956 and 1967.”

1RXS J232953.9+062814 J. Biggs, Perth Observatory, reports that four exposures were taken of

this object (cf. IAUC 7747) by Biggs and S. Walsh with the Mike Candy Telescope on Nov. 8.62, yielding magnitude R = 14.8 ± 0.5 and the following position: α = 23^{h}29^{m}54^{s}.2, δ = +6^{o}28′12′′ (equinox 2000.0); the uncertainty of ± 1′′ is within the 11′′ uncertainty of its x-ray position (cf. Voges et al. 1999, A.Ap. 349, 389). Biggs adds that a star in the USNO-A2.0 catalogue at R = 15.6 has position end figures 54^{s}.36, 10′′.4.

2001 November 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green